930. (24 pts) Page 44. Problem 30: a, b,c. d. 30. The following data give the weekly amounts spent on groceries for a sample of 45 households $271 $33 $159 76 $227 $337 $295 $319 $250 162 232 303 41 335 170 188 320 279 192 116 429 205 181 100 294 279 266 321 151 240 474 297 570 342 279 235 434 123 325 199 246 309 278 50 a. How many classes would you recommend? b. What class interval would you suggest? c. What would you recommend as the lower limit of the first class? d. Organize the data into a frequency distribution.
Answer #1

a) There are 45 groups,

2^5=32

2^6=64

32<45<64

I would recommend 6 classes

b) Range = 570-41=529

529/6 round to 82 I would suggest 90

c) 40

d) Organize the information into a frequency distribution.

Class Interval Frequency
40 - 130 6
130 -210 10
210 -300 15
300 - 390 10
390 - 480 3
480 -570 1
Total 45